OpenStudy (anonymous):

The line that passes through (1, 5) and (-2, 3) in standard form

OpenStudy (texaschic101):

Lets find the slope first by using the slope formula.. slope(m) = (y2 - y1) / (x2 - x1) (1,5)...x1 = 1 and y1 = 5 (-2,3)...x2 = -2 and y2 = 3 now we sub slope(m) = (3 - 5) / (-2 - 1) slope(m) = -2/-3 slope(m) = 2/3 now that we have the slope(m), we need to find the y intercept(b) y = mx + b slope(m) = 2/3 you can use either of your points.. (1,5)...x = 1 and y = 5 now we sub 5 = 2/3(1) + b 5 = 2/3 + b 5 - 2/3 = b 15/3 - 2/3 = b 13/3 = b so now we have : y = 2/3x + 13/3...however, we need it in standard form. Ax + By = C y = 2/3x + 13/3 -- multiply entire equation by 3 3y = 2x + 13 -- subtract 2x from both sides -2x + 3y = 13 -- multiply by -1 to make x positive 2x - 3y = -13 <== standard form

OpenStudy (texaschic101):

any questions ?